OpenStudy (anonymous):

Solve. p divided by 8 equals twelve p equals two-thirds p equals 1 and one-half p equals 4 p equals ninety-six @ivettef365

OpenStudy (ivettef365):

p/8= 12 just multiply 8 x 12

OpenStudy (texaschic101):

p/8 = 12 (multiply by 8) (8)p/8 = (8)12 p = 96